After the first term, each term in a sequence is 3 greater than 1/3 of the preceeding term. If t is the first term of the sequence and t (cannot be = to) 0, what is the ratio of the second term to the first term?
(A) t+9/3
(B) t+3/3
(C) t+9/3t
(D) t+3/3t
(E) 9-2t/3

Answer by stanbon(75887)

After the first term, each term in a sequence is 3 greater than 1/3 of the preceeding term. If t is the first term of the sequence and t (cannot be = to) 0, what is the ratio of the second term to the first term?
-----------
1st term = "t"
2nd term = (1/3)t + 3
-----------------
Ratio = [(1/3)t+3]/t = (1/3) + (3/t) = (t+9)/3t
=============
Cheers,
Stan H.